The Paulson Company's year-end balance sheet is shown below. Its cost of common equity is 16%, its before-tax cost of debt is 8%, and its marginal tax rate is 25%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,162. The firm has 576 shares of common stock outstanding that sell for $4.00 per share

Calculate Paulson's WACC using market-value weights. Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Cost of Equity = 16%

Cost of Debt after tax = 8% * (1-0.25) = 6%

Total Capital = Debt + Equity

= 1162 + 576*4

= 1162 + 2304

= 3466

WACC = Cost of Equity * Weight of Equity + Cost of Debt after tax * Weight of Debt

= 16% * 2304/3466 + 6% * 1162/3466

= 12.65%

WACC is 12.66%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.